# Detection and Reconstruction of a Random Hypergraph from Noisy Graph Projection

### v1

@@ 1.2 The reconstruction threshold
=\ = #$\#\Big{\{}A:|E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\Big{\}}$
[…]
@@
=\ = $\sum_{\begin{subarray}{c}I\geq\iota s\binom{n}{d}\\ J\geq 0\end{subarray}}\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime}:|\mathcal{H}\cap\mathcal{H}^{\prime}|=I\\ |E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|=\binom{d}{2}I+J\end{subarray}}\mu^{\prime}(\mathcal{H})\mu^{\prime}(\mathcal{H}^{\prime})\sum_{\begin{subarray}{c}A:\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H}^{\prime})\subset A\\ |E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\end{subarray}}\mathbb{P}^{\prime}(A\mid\mathcal{H})\mathbb{P}^{\prime}(A\mid\mathcal{H}^{\prime})\,.$ (5.10)
@@
#$\#\Big{\{}A:\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H}^{\prime})\subset A,|E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\Big{\}}$
@@
=\ = $[1+o(1)]\cdot\mathbb{P}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Big{(}|\mathcal{H}\cap\mathcal{H}^{\prime}|=I,|E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|=I\tbinom{d}{2}+J\Big{)}\,.$ (5.12)
@@
$\mathbb{P}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Big{(}|\mathcal{H}\cap\mathcal{H}^{\prime}|=I\Big{)}\leq$ $\binom{\binom{n}{d}}{(1+o(1))s\binom{n}{d}-I}\binom{(1+o(1))s\binom{n}{d}}{I}\Bigg{/}\binom{\binom{n}{d}}{(1+o(1))s\binom{n}{d}}$
@@
$\mathbb{P}_{(\mathcal{K},\mathcal{K}^{\prime})\sim\widetilde{\mu}\otimes\widetilde{\mu}}\Big{(}\mathcal{K}\cap\mathcal{K}^{\prime}=\emptyset\Big{)}=(1-s^{2})^{\binom{n}{d}}\geq\exp\big{(}-s^{2}\tbinom{d}{2}\big{)}=\exp\big{(}o(1)\cdot I\big{)}\,,$
@@
To see this, note that for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$, we have that $(\widehat{\mathcal{H}},\mathcal{H})$ and $(\widehat{\mathcal{H}},\mathcal{H}^{\prime})$ are equal in law. Thus, we have
$o(1)=\mathbb{P}\Big{(}|\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\tbinom{n}{d}\Big{)}$ (|^$\geq\mathbb{P}\Big{(}|\mathcal{H}^{\prime}\triangle\widehat{\mathcal{H}}|\leq o(1)\cdot s\tbinom{n}{d}\Big{)}+\mathbb{P}\Big{(}|\mathcal{H}\triangle\widehat{\mathcal{H}}|\leq o(1)\cdot s\tbinom{n}{d}\Big{)}$
$=2\mathbb{P}\Big{(}|\mathcal{H}\triangle\widehat{\mathcal{H}}|\leq o(1)\cdot s\tbinom{n}{d}\Big{)}\,,$
where the inequality follows from the fact that
$|\mathcal{H}\triangle\widehat{\mathcal{H}}|+|\mathcal{H}^{\prime}\triangle\widehat{\mathcal{H}}|\geq|\mathcal{H}\triangle\mathcal{H}^{\prime}|\,.$
This yields Proposition 1.9. We now prove (6.1). For any $\mathcal{H}\in\{0,1\}^{\binom{[n]}{d}}$ and $A\in\{0,1\}^{\binom{[n]}{2}}$, define
@@
[…]
@@
[…]
@@
#$\#\mathsf{Blur}(\mathcal{H};A)\geq\binom{\binom{n}{d}-(1+o(1))s\binom{n}{d}}{|\mathsf{Empty}(\mathcal{H};A)|}\geq\exp\Big{(}\Omega(1)\cdot s\tbinom{n}{d}\log n\Big{)}\,.$
@@
=\ = $|E(\mathcal{P}(\mathsf{Empty}(\mathcal{H};A)))|-|E(\mathcal{P}(\mathcal{H}^{\prime}\cap\mathsf{Empty}(\mathcal{H};A)))\setminus E(A)|$
[…]
@@
[…]
@@
#$\#\Big{\{}\mathcal{H}:|\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\tbinom{n}{d}\Big{\}}\leq 2^{|\mathcal{H}^{\prime}|}\cdot\binom{\binom{n}{d}}{o(1)\cdot s\binom{n}{d}}=\exp\Big{(}o(1)\cdot s\tbinom{n}{d}\log n\Big{)}\,.$
@@ Acknowledgment
S. Gong and Z. Li are partially supported by National Key R&\&&D Program of China (Project No. 2023YFA1010103) and NSFC Key Program (Project No.12231002).
@@ Appendix A Detection with a modified null distribution
In the main text, we have cited the presence of a detection-recovery gap as a source of difficulty for obtaining tight thresholds for recovery. One might wonder whether this can be remedied simply by choosing a better null distribution $\mathbb{Q}$ that is harder to distinguish from the planted distribution $\mathbb{P}$. In this section, we will argue that this approach does not work. Recall Definition 1.1. Define ~=~$\widetilde{\mathbb{Q}}=\widetilde{\mathbb{Q}}_{n}$ to be the Erdős-Rényi graph with edge-density
@@
~$\widetilde{\mathbb{Q}}\Big{(}g(A)<q^{\binom{d}{2}}\tbinom{n}{d}+\tfrac{1}{2}ps\tbinom{n}{d}\Big{)}=1-o(1)\,.$
@@
Denote $\kappa=\max\{\delta,\beta\}$ and recall from (A.1) that $\widetilde{q}=n^{-1+\kappa+o(1)}$. When $\kappa<\tfrac{d-3}{d-1}$, it can be easily checked using Markov’s inequality that ~$\widetilde{\mathbb{Q}}(g(A)=0)=1-o(1)$. When $\kappa\geq\tfrac{d-3}{d-1}$, we have
@@
Note that for $2\leq m\leq d$, we have #$\#\{\Psi,\Phi:|V(\Psi)\cap V(\Phi)|=m\}=n^{2d-m+o(1)}$ and for $|V(\Psi)\cap V(\Phi)|=m$ we have
@@
Note that when $\kappa>\tfrac{d-3}{d-1}$, we have $-m+(1-\kappa)\binom{m}{2}$ is monotone decreasing in $m$ when $2\leq m\leq d$. Thus,
$\operatorname{Var}_{\widetilde{\mathbb{Q}}}(g(A))=\mathbb{E}_{\widetilde{\mathbb{Q}}}[g(A)]^{2}\cdot n^{-1-\kappa+o(1)}\,.$
@@
$\operatorname{Var}_{\widetilde{\mathbb{Q}}}(g(A))^{\tfrac{1}{2}}\leq n^{d-\binom{d}{2}(1-\kappa)-\tfrac{1+\kappa}{2}+o(1)}\leq n^{1+\delta-\Omega(1)}\,,$
where the last inequality follows from our assumption that
$d-\tbinom{d}{2}(1-\kappa)-\tfrac{1+\kappa}{2}=1+\delta-\Omega(1)\mbox{ when }\beta,\delta<\tfrac{d-1}{d+1}\,.$
@@
Note that $\widetilde{q}-q=n^{-1+\delta+o(1)}$ and $q=n^{-1+\beta+o(1)}$, thus we have for all $10\leq k\leq\tbinom{d}{2}-$ 1
@@
~$\widetilde{\mathbb{Q}}\Big{(}g(A)<q^{\binom{d}{2}}\tbinom{n}{d}+\tfrac{1}{2}ps\tbinom{n}{d}\Big{)}=1-o(1)\,.$
In addition, note that under $\mathbb{P}$ we have $|\mathcal{H}|\geq 0.9ps\tbinom{n}{d}$ with probability $1-o(1)$. Thus, it is clear that

### v2

@@ 2.2 The reconstruction threshold
$=\$$\#\Big\{A:|E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\Big\}$
$\leq\$$\binom{\binom{n}{2}}{q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}}\leq\exp\Big(\big(q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\big)\log(q^{-1})\Big)\,$, (6.9)
@@
[…]
@@
$\#\Big\{A:\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H}^{\prime})\subset A,|E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\Big\}$
@@
$=\$$[1+o(1)]\cdot\mathbb{P}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu^{\prime}\otimes\mu^{\prime}}\Big(|\mathcal{H}\cap\mathcal{H}^{\prime}|=I,|E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|=I\tbinom{d}{2}+J\Big)\,.$ (6.12)
@@
$\mathbb{P}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu^{\prime}\otimes\mu^{\prime}}\Big(|\mathcal{H}\cap\mathcal{H}^{\prime}|=I\Big)\leq\$$\binom{\binom{n}{d}}{(1+o(1))s\binom{n}{d}-I}\binom{(1+o(1))s\binom{n}{d}}{I}\Bigg/\binom{\binom{n}{d}}{(1+o(1))s\binom{n}{d}}$
@@
$\mathbb{P}_{(\mathcal{K},\mathcal{K}^{\prime})\sim\widetilde{\mu}\otimes\widetilde{\mu}}\Big(\mathcal{K}\cap\mathcal{K}^{\prime}=\emptyset\Big)=(1-s^{2})^{\binom{n}{d}}\geq\exp\big(-s^{2}\tbinom{n}{d}\big)=\exp\big(-o(1)\cdot I\big)\,$,
@@
To see this, note that for any estimator $\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$, we have that (^,)($\widehat{\mathcal{H}}$$\mathcal{H}$) and (^,)($\widehat{\mathcal{H}}$$\mathcal{H}^{\prime}$) are equal in law. Combining the fact that
[…]
we then have
[…]
where the first inequality follows from (7.1). This yields Proposition 2.9.
We now prove (7.1). For any $\mathcal{H}\in\{0,1\}^{\binom{[n]}{d}}$ and $A\in\{0,1\}^{\binom{[n]}{2}}$, define
@@
$\mathcal{G}$$\Big\{$ $(\mathcal{H},A):|\mathcal{H}|=(1+o(1))s\tbinom{n}{d},\mathsf{Empty}(\mathcal{H};A)\geq\Omega(1)\cdot s\tbinom{n}{d}$; (7.3)
@@
$\mathsf{Blur}(\mathcal{H};A)\triangleq\Big\{$ $\mathcal{H}^{\prime}\in\{0,1\}^{\binom{[n]}{d}}:|\mathcal{H}^{\prime}|=|\mathcal{H}|,\mathcal{H}\setminus\mathsf{Empty}(\mathcal{H};A)\subset\mathcal{H}^{\prime}\Big\}\,.$ (7.4)
@@
$\#\mathsf{Blur}(\mathcal{H};A)\geq\binom{\binom{n}{d}-(1+o(1))s\binom{n}{d}}{|\mathsf{Empty}(\mathcal{H};A)|}\geq\exp\Big(\Omega(1)\cdot s\tbinom{n}{d}\log n\Big)\,.$
@@
$=\$$|E(\mathcal{P}(\mathsf{Empty}(\mathcal{H};A)))|-|E(\mathcal{P}(\mathcal{H}^{\prime}\cap\mathsf{Empty}(\mathcal{H};A)))\setminus E(A)|$
$\geq\$$(1-o(1))\tbinom{d}{2}|\mathsf{Empty}(\mathcal{H};A)|-\tbinom{d}{2}|\mathsf{Empty}(\mathcal{H};A)|=-o(1)\cdot s\tbinom{n}{d}\,$.
@@
$\leq\$$\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime};A:(\mathcal{H},A),(\mathcal{H}^{\prime},A)\in\mathcal{G}\\ |\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\binom{n}{d}\end{subarray}}\frac{\mathbb{P}(\mathcal{H},A)\mathbb{P}(\mathcal{H}^{\prime},A)}{\mathbb{P}(A)}+2\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime};A\\ (\mathcal{H},A)\not\in\mathcal{G}\end{subarray}}\frac{\mathbb{P}(\mathcal{H},A)\mathbb{P}(\mathcal{H}^{\prime},A)}{\mathbb{P}(A)}$ (7.6)
@@
$\#\Big\{\mathcal{H}:|\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\tbinom{n}{d}\Big\}\leq 2^{|\mathcal{H}^{\prime}|}\cdot\binom{\binom{n}{d}}{o(1)\cdot s\binom{n}{d}}=\exp\Big(o(1)\cdot s\tbinom{n}{d}\log n\Big)\,.$
@@ Acknowledgment
The main part of the work was carried out when Q. Xu was at Peking University. S. Gong and Z. Li are partially supported by National Key R&\&D Program of China (Project No. 2023YFA1010103) and NSFC Key Program (Project No.12231002).
@@ Appendix A Detection with a modified null distribution
In the main text, we have cited the presence of a detection-recovery gap as a source of difficulty for obtaining tight thresholds for recovery. One might wonder whether this can be remedied simply by choosing a better null distribution $\mathbb{Q}$ that is harder to distinguish from the planted distribution $\mathbb{P}$. In this section, we will argue that this approach does not work. Recall Definition 2.1. Define $\widetilde{\mathbb{Q}}=\widetilde{\mathbb{Q}}_{n}$ to be the Erdős-Rényi graph with edge-density
@@
$\widetilde{\mathbb{Q}}\Big(g(A)<q^{\binom{d}{2}}\tbinom{n}{d}+\tfrac{1}{2}ps\tbinom{n}{d}\Big)=1-o(1)\,.$
@@
Denote $\kappa=\max\{\delta,\beta\}$ and recall from (A.1) that $\widetilde{q}=n^{-1+\kappa+o(1)}$. When $\kappa<\tfrac{d-3}{d-1}$, it can be easily checked using Markov’s inequality that $\widetilde{\mathbb{Q}}(g(A)=0)=1-o(1)$. When $\kappa\geq\tfrac{d-3}{d-1}$, we have
@@
Note that for $2\leq m\leq d$, we have $\{\Psi,\Phi:|V(\Psi)\cap V(\Phi)|=m\}=n^{2d-m+o(1)}$ and for $|V(\Psi)\cap V(\Phi)|=m$ we have
@@
Note that when $\kappa>\tfrac{d-3}{d-1}$, we have the maximum of $m+(1-\kappa)\binom{m}{2}$ where $2\leq m\leq d$ is taken by either $m=2$ or $m=d$. Thus,
$\operatorname{Var}_{\widetilde{\mathbb{Q}}}(g(A))=\mathbb{E}_{\widetilde{\mathbb{Q}}}[g(A)]^{2}\cdot n^{-\min\{(1+\kappa),d-\binom{d}{2}(1-\kappa)\}+o(1)}\,.$
@@
$\operatorname{Var}_{\widetilde{\mathbb{Q}}}(g(A))^{\tfrac{1}{2}}\leq n^{d-\binom{d}{2}(1-\kappa)-\tfrac{1}{2}\min\{(1+\kappa),d-\binom{d}{2}(1-\kappa)\}+o(1)}\leq n^{1+\delta-\Omega(1)}\,$,
where the last inequality follows from our assumption that (recall $\kappa=\max\{\delta,\beta\}$)
$\tfrac{d-\tbinom{d}{2}(1-\kappa)}{2},\ d-\tbinom{d}{2}(1-\kappa)-\tfrac{1+\kappa}{2}=1-\Omega(1)\mbox{ when }\beta,\delta<\tfrac{d-1}{d+1}\,.$
@@
Note that $\widetilde{q}-q=n^{-1+\delta+o(1)}$ and $q=n^{-1+\beta+o(1)}$, thus we have for all $0\leq k\leq\tbinom{d}{2}-1$
@@
$\widetilde{\mathbb{Q}}\Big(g(A)<q^{\binom{d}{2}}\tbinom{n}{d}+\tfrac{1}{2}ps\tbinom{n}{d}\Big)=1-o(1)\,.$
In addition, note that under $\mathbb{P}$ we have $|\mathcal{H}|\geq 0.9ps\tbinom{n}{d}\gg q^{\binom{d}{2}}\tbinom{n}{d}$ with probability 1o(1)1-o(1) when $\beta<\tfrac{d-1}{d+1}$. Thus, it is clear that